OpenStudy (anonymous):

A solution of 56% fertilizer is to be mixed with a solution of 25% fertilizer to form 155 L of a 51% solution. How many liters of the 56% solution must be used?

OpenStudy (anonymous):

To create a breakfast beverage, cherry juice in two concentrations, 19% and 60%, must be combined into a solution that will be mixed with another type of juice to produce the beverage. If 70 gallons of 19% juice is used, how many gallons of the 60% juice must be used to obtain a 50% cherry juice solution?

OpenStudy (dumbcow):

 From this table we get 2 equations: \[x + y = 155\] \[56x + 25y = 51(155)\] solve with substitution --> y = 155 - x plug into 2nd equation --> 56x + 25(155-x) = 51(155) solve for x

OpenStudy (dumbcow):

*sorry i switched the 51 and 56 in the table

OpenStudy (anonymous):

hmm I think I understand it

OpenStudy (anonymous):

so x= 130 I solved for x

OpenStudy (dumbcow):

ok good and you can use this table for any of these type of mixing questions 

OpenStudy (dumbcow):

yes x = 130 good job